OpenStudy (anonymous):

a ball is shot straight up from the surface of the earth with an initial speed of 19.6 m/s. Neglect any effects due to air resistance. 29. What is the magnitude of the ball's displacement from the starting point after 1.00 has elapsed? 30. what maximum height will the ball reach? 31. how much time elapses between the throwing of the ball and its return to the original launch point?

OpenStudy (anonymous):

voy=19.6 cause we are dealing with y, right?

OpenStudy (anonymous):

1.00 second has elapsed for problem 29

OpenStudy (anonymous):

for number 19 is it seconds? 1.00seconds? and yes, we're dealing with y

OpenStudy (anonymous):

oh okay, i just asked that XD

OpenStudy (anonymous):

my vault lol

OpenStudy (anonymous):

so the magnitude is v y final?

OpenStudy (anonymous):

oh no, it is 14.7

OpenStudy (anonymous):

where did I go wrong

OpenStudy (anonymous):

unless if magnitude is voy-gt/2

OpenStudy (anonymous):

though I am not sure

OpenStudy (anonymous):

do you want to solve it here or i can just solve it on my own and just tell me what you got.? we don't need to find for vfy for number 29, just use d=viyt+1/2at^2 viy (initial velocity is 19.6 m/s)

OpenStudy (anonymous):

no gravity either?

OpenStudy (anonymous):

acceleration=gravity

OpenStudy (anonymous):

so a=9.8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if I am not mistaken

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

so what did you get for 29?

OpenStudy (anonymous):

I get d=19.6t+1/2at squared

OpenStudy (anonymous):

a=9.8 right?

OpenStudy (anonymous):

19.6+24.01=43.61 but that isn't the correct answer

OpenStudy (anonymous):

remeber t=1.00 sec and yes a=9.8m/s^2 and vi=19.6 m/s

OpenStudy (anonymous):

the correct answer is 14.7

OpenStudy (anonymous):

so I am messing up somewhere

OpenStudy (anonymous):

ahh 4.9 and 19.6

OpenStudy (anonymous):

I made a carless mistake

OpenStudy (anonymous):

so 19.6-4.9=14.7

OpenStudy (anonymous):

because velocity is up and acceleration due to gravity is down so it will be 19.6-4.9=14.7

OpenStudy (anonymous):

and you're right again XD i guess i need to practice typing fast O.O

OpenStudy (anonymous):

ahh I get it

OpenStudy (anonymous):

I learn the most when I practice with others than practicing alone personally

OpenStudy (anonymous):

what maximum height will the ball reach?

OpenStudy (anonymous):

so Vy=Voy-2g(y-yo)

OpenStudy (anonymous):

opps, thats not wright...

OpenStudy (anonymous):

let's list what we have so far a= -9.8 m/s^2 vi= 19.6 m/s so for number 30, Vf will be 0 m/s for easier calculation, use \(\Large V_f^2=V_i^2 + 2ad\)

OpenStudy (anonymous):

30 is 19.6 hmm...

OpenStudy (anonymous):

what a minute, the ball was shot in the air for 1 second

OpenStudy (anonymous):

19.6 times 1=19.6

OpenStudy (anonymous):

so the maximum height the ball will reach is 19.6 at 1 second, right?

OpenStudy (anonymous):

or am I approaching it the wrong way?

OpenStudy (anonymous):

we're not sure if the ball reached the maximum height in 1 second, so it's better not to use the time

OpenStudy (anonymous):

for number 29, they just asked where will the ball after 1 sec..

OpenStudy (anonymous):

so the ball will reach a maximum height

OpenStudy (anonymous):

of 19.6

OpenStudy (anonymous):

that's the initial velocity

OpenStudy (anonymous):

the maximum height is apparently 19.6m as well according to the answer key

OpenStudy (anonymous):

maybe it is just coincidence?

OpenStudy (anonymous):

anyways

OpenStudy (anonymous):

we got to wrap this up soon

OpenStudy (anonymous):

how much time elapses between the throwing of the ball and its return to the original launch point?

OpenStudy (anonymous):

oh you already solved it , wait give me a sec, i'll solve it too , sorry :/

OpenStudy (anonymous):

its fine

OpenStudy (anonymous):

yeah, i got the same thing 19.6 okay next question XD how much time elapses between the throwing of the ball and its return to the original launch point? d=0 so we can use d=vit + 1/2 gt^2 now we need to use the quadratic formula ^.^

OpenStudy (anonymous):

still there? O.o

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so 4.9t squared+19.6t-14.7=0?

OpenStudy (anonymous):

don't forget the direction the gravity is down, velocity is up set up as positive, gravity will be negative

OpenStudy (anonymous):

-4.9tsquared+19.6t-14.7?

OpenStudy (anonymous):

where did you get 14.7? it's the maximum height right? we actually dont need it because it's asking for the time when the object came back to its original position so it will be zero: \(-4.9t^2+19.6t+0\)

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

derp

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

thank so much for all the help

OpenStudy (anonymous):

9: 57 pm here 3 mins more , anything else?? ^.^ good luck on your exams!! you can do it!! remember your directions (positive or negative), your units

OpenStudy (anonymous):

when do you know when something is positive and when something is negative?

OpenStudy (anonymous):

it depends on which direction you set as positive or negative

OpenStudy (anonymous):

shot a rocket up

OpenStudy (anonymous):

between shooting a rocket down

OpenStudy (anonymous):

if you set up as positive , down will be negative and vice versa for objects going up, it's more efficient to use \(up\) as positive for objects thrown downwards, it's more efficient to use \(down\) as positive

OpenStudy (anonymous):

thank man, I really do appreciate it

OpenStudy (anonymous):

no problem (^_^) i'm doing physics too ! so no worries ;p